Given the function y=x^2 + 9x + 30, what is the least possible value of y for integer values of x?
Answer by Onmyoji.S(32)

Given the function y=x^2 + 9x + 30, what is the least possible value of y for integer values of x?
_________❤________________❤_________________❤______________
x=-9/2=-4.5
y=-4.5^2+9*4.5+30
y=50.25
_________❤________________❤_________________❤______________
[[X|Y]
[01|41
[02|52
[03|66
[00|30
[-1|22
[-2|16
[-3|21
❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤♡❤
———————♥——————————//(◕‿‿◕)\\—————————♥———————
I Guess.